OpenStudy (anonymous):

Solve for X & check all that apply. 2700=300*2^10x A. x=ln(9)/10 B. x=ln(9)/10*ln(2) C. x=log(9)/10 D. x=log(9)/10*log(2)

OpenStudy (anonymous):

is that 2700=300*2^(10x) ? if so then steps would be: 2700/300 = 2^(10x) then take the Ln of both sides: Ln(9) = Ln(2)^(10x) Now you can drop the 10x: Ln(9) = 10x*Ln(2) then just move the 10*Ln(2) to the other side: Ln(9)/10*Ln(2)=x

OpenStudy (anonymous):

So B would be the only possible option and none of the other ones would apply?

OpenStudy (anonymous):

D would be correct also because of the change of base formula

OpenStudy (anonymous):

oh okay so just B & D?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

thank you so much! :)

OpenStudy (anonymous):

You're welcome!